# Bulk-Boundary Correspondence in 2D Topological Photonics: Analysis and Simulation

### v2

@@ VI Acknowledgment
The 2D analysis consists of three main steps: 1) differentiate the wave equation (2) with respect to $s=k^{2}=(\omega/c)^{2}$, leading to an equation for the derivative $g=\partial_{s}e$; 2) inner-multiply this equation for $g$ with $e$ and integrate by parts, to arrive at the boundary values of the fields; 3) algebraically simplify the result to obtain a relationship for the DtN operator. For 1D problems, this plan was pioneered in [20], with some simplification given in [24].
In 1D, the boundary admittance (inverse of impedance) is a scalar; in 2D, it turns into the Dirichlet-to-Neumann (DtN) operator $\mathcal{D}(s)$:
@@
where the inner products over the strip $S$ and its boundary $n=0$ are defined as follows:
@@
The last term can be transformed due to (24):

### v3

@@ VI Acknowledgment
The 2D analysis consists of two main steps: 1) differentiate the wave equation (2) with respect to $s=k^{2}=(\omega/c)^{2}$, leading to an equation for the derivative $g=\partial_{s}e$; 2) inner-multiply this equation for $g$ with $e$ and integrate by parts, arriving at the boundary values of the fields and at a relationship for the DtN operator. For 1D problems, this plan was pioneered in [20] and slightly simplified in [24]. In 1D, the boundary admittance (inverse of impedance) is a scalar; in 2D, it turns into the Dirichlet-to-Neumann (DtN) operator $\mathcal{D}(s)$ (for brevity, dependence on $\beta$ is not explicitly indicated):
@@
The inner products over the strip $S$ and its boundary $n=0$ are defined as follows:
@@
The last term can be transformed due to (24), and we have